OpenStudy (eric_d):

Simple Matrix 4)a 4)C

OpenStudy (eric_d):

http://prntscr.com/4bq00a

OpenStudy (anonymous):

These are just simple ones..

OpenStudy (anonymous):

See, here I am giving you just general example: \[\left[\begin{matrix}a & b \\ c & d\end{matrix}\right]\left[\begin{matrix}e & f \\ g & h\end{matrix}\right]\implies \left[\begin{matrix}ae + bg & ce + cg \\ af + bh & af + dh\end{matrix}\right]\]

OpenStudy (anonymous):

Oh, thanks @hartnn you are here..

OpenStudy (anonymous):

I have to go,it is urgent, just help @eric_d and when I will come online next, I will carry it on.. :)

OpenStudy (anonymous):

There you are given with matrix A and B.. Just substitute the values in place of a,b,c,d,e,f,g,h and try to solve as I have done it above..

OpenStudy (eric_d):

ok

OpenStudy (eric_d):

what about 4)c)

hartnn (hartnn):

both 4a and 4c are of same type learn one, other is similar http://www.mathsisfun.com/algebra/matrix-multiplying.html

OpenStudy (anonymous):

It is about addition first and then multiplication. You can add corresponding elements of matrix..

OpenStudy (anonymous):

\[\left[\begin{matrix}a & b \\ c & d\end{matrix}\right] + \left[\begin{matrix}e & f \\ g & h\end{matrix}\right]\implies \left[\begin{matrix}a +e & b+f \\ c + g & d + h\end{matrix}\right]\]

OpenStudy (eric_d):

Thanks @waterineyes for the general examples.. and also @hartnn